OpenStudy (bekkah323):

find an equation of the circle whose center has coordinates (-1,2), with radius 3

OpenStudy (anonymous):

Equation of a circle is \[(x - h)^2 + (y - k)^2 = r^2\] where (h,k) is the center and r is the radius therefore the equation is \[(x - (-1))^2 + (y - 2)^2 = 3^2\] simplify and you get \[(x + 1)^2 + (y-2)^2 = 9\]